Grade 11Mechanics

body starting fron rest has an accleration of 20 m/s² calculate the distance travelled by it in 6 th sec

Sanchit
9 Years agoGrade 11

1 Answer

Satyam Garg
9 Years ago
Using formulaS(n) = u + a/2(2n-1)Since the body has started from rest its u=0Substituting values, we getS(6) = 0 + 20/2(2*12-1)Therefore S(6) = 130mHence distance travelled in 6th second is 130m